OpenStudy (anonymous):

A=1 b=3 can someone help me with this please I can figure out the answer A=1 b=3 (1+3)^2=1^2+3^2 4^2=3+6

OpenStudy (anonymous):

so it would be (1+3)^2 I already got that

OpenStudy (anonymous):

I cant figure the answer out

OpenStudy (anonymous):

no I am trying to figure out the answer for A=1 b=3 (1+3)^2=1^2+3^2 4^2=3+6